OpenStudy (anonymous):

Find the domain, range and asymptote of y=10-e^(-x)

OpenStudy (anonymous):

missing something ?

OpenStudy (anonymous):

whatever it is suppose to be make sure what is inside the log is larger than zero that is all

OpenStudy (anonymous):

One sec. Wrong function

OpenStudy (anonymous):

\[-x>0\] solve for \(x\)

OpenStudy (anonymous):

So x is greater than 0? Domain is x element of the real numbers?

OpenStudy (anonymous):

@satellite73

OpenStudy (rational):

domain is all the x values that you can plug into a function

OpenStudy (rational):

look at the given expression, are there any x values that make the expression undefined ?

OpenStudy (anonymous):

no?

OpenStudy (rational):

so the domain is all real numbers : \((-\infty,~ +\infty)\)

OpenStudy (anonymous):

Okay, how do you get the range?

OpenStudy (rational):

find the limits as \(x\to+\infty\) and \(x\to -\infty\)

OpenStudy (anonymous):

The range is the real numbers?

OpenStudy (rational):

how ? what did u get for the limits ?

OpenStudy (anonymous):

I wasn't really sure if I got the limits... but it looks like it goes infinity and neg. infinity..

OpenStudy (rational):

"it looks like" is a forbidden phrase in calculus, work the limits and you will be surprised im sure

OpenStudy (anonymous):

Oh... Can you please help with one of the limits? Just to make sure I'm doing it correctly..

OpenStudy (rational):

\[\large{\begin{align} \lim\limits_{x\to+\infty} 10-e^{-x} &= \lim\limits_{x\to+\infty} 10- \lim\limits_{x\to+\infty} e^{-x}\\~\\ &=\lim\limits_{x\to+\infty} 10- e^{\lim\limits_{x\to+\infty}-x}\\~\\ &=10- e^{-\infty}~~*\\~\\ &=10-0\\~\\ &=10\\~\\ \end{align}}\]

OpenStudy (rational):

so the graph of given function flattens to 10 as you walk extreme right on x axis

OpenStudy (rational):

work the other limit to find out what happens to the function as you go extreme left

OpenStudy (anonymous):

how did e^-infinity turns out to be a zero?

OpenStudy (rational):

good question, i have used exponent property : \[\large a^{-n} = \frac{1}{a^n}\]

OpenStudy (rational):

\[\large e^{-\infty} = \frac{1}{e^{\infty}} = \frac{1}{\text{think of very big positive quantity here}} = 0\]

OpenStudy (anonymous):

oh I see.

OpenStudy (anonymous):

It approaches to zero

OpenStudy (anonymous):

So the other limit goes to infinity.

OpenStudy (rational):

sure ?

OpenStudy (rational):

does it go to "positive" infinity or "negative" infinity

OpenStudy (anonymous):

negative infinity

OpenStudy (rational):

good, can you guess the range now keep in mind 10-e^(-x) is a continuous function

OpenStudy (rational):

also it is a strictly increasing function work the first derivative quick if you want

OpenStudy (rational):

remember how to prove if a function is increasing ?

OpenStudy (anonymous):

Not really...

OpenStudy (anonymous):

But is the range this (-infinity,10)

OpenStudy (rational):

In \([a,b]\), if \(f'(x)\) is positive, then \(f(x)\) is a strictly increasing in that interval.

OpenStudy (rational):

yes range is (-infinity, 10)

OpenStudy (rational):



OpenStudy (anonymous):

So there's horizontal asymptote on y=10?

OpenStudy (rational):

Yes, also thats clear from graph

OpenStudy (anonymous):

I think I got it from here. Thank you for your help! :)

OpenStudy (rational):

yw